# Standard Gaussian Process Can Be Excellent for High-Dimensional Bayesian Optimization

### v3

@@ Abstract
There has been a long-standing and widespread belief that Bayesian Optimization (BO) with standard Gaussian process (GP), referred to as standard BO, is ineffective in high-dimensional optimization problems. While this belief sounds reasonable, strong empirical evidence is lacking. In this paper, we systematically investigated BO with standard GP regression across a variety of synthetic and real-world benchmark problems for high-dimensional optimization. We found that, surprisingly, when using Matérn kernels and Upper Confidence Bound (UCB), standard BO consistently achieves top-tier performance, often outperforming other BO methods specifically designed for high-dimensional optimization. Contrary to the stereotype, we found that standard GP equipped with Matérn kernels can serve as a capable surrogate for learning high-dimensional functions. Without strong structural assumptions, BO with standard GP not only excels in high-dimensional optimization but also is robust in accommodating various structures within target functions. Furthermore, with standard GP, achieving promising optimization performance is possible via maximum a posterior (MAP) estimation with diffuse priors or merely maximum likelihood estimation, eliminating the need for expensive Markov-Chain Monte Carlo (MCMC) sampling that might be required by more complex surrogate models. In parallel, we also investigated and analyzed alternative popular settings in running standard BO, which, however, often fail in high-dimensional optimization. This might link to the a few failure cases reported in literature. We thus advocate for a re-evaluation and in-depth study of the potential of standard BO in addressing high-dimensional problems. The code for our experiment is available at https://github.com/XZT008/Standard-GP-is-all-you-need-for-HDBO.
@@ 1 Introduction
Despite numerous successes, there has been a widespread belief that BO with standard GP regression, referred to as standard BO, is limited to low-dimensional optimization problems [8, 25, 6, 24, 18, 32, 19]. It is commonly thought that the number of optimization variables should not exceed 15 or 20, as beyond this threshold, BO is prone to failure [8, 25]. This belief may partly stem from the intuition that, as a kernel method, GPs struggle to handle high-dimensional inputs for covariance modeling, rending them incapable of providing an expressive prior for high-dimensional functions.
Kernel GP Training Acquisition Function
Matérn ✓ L-BFGS/MAP/Diffuse prior ✓ UCB ✓
ADAM/MAP/Diffuse prior ✓
L-BFGS/MAP/Concentrated prior ✗ EI ✗
SE ✗ ADAM/MAP/Concentrated prior ✗
L-BFGS/MLE ✗ TS ✗
ADAM/MLE ✓
Table 1: Favorable and unfavorable settings of running standard BO for high-dimensional optimization. SE: square exponential; EI: expected improvement; TS: Thompson sampling; MLE: maximum likelihood estimation.
@@
A continuous line of research is dedicated to developing novel high-dimensional BO methods. The fundamental strategy is to impose strong structural assumptions into surrogate modeling so as to avoid directly dealing with high-dimensional inputs. One class of methods assumes a decomposition structure within the functional space [16, 26, 10, 34]. In other words, the target function can be expressed as the summation of a group of low-dimensional functions, each operating over a small number of variables. Another family of methods assumes that the inputs are intrinsically low-rank. These methods either project the original inputs into a low-dimensional space [32, 25, 18] or use sparse-inducing priors to trim down massive input variables [6]. Subsequently, GP surrogates are built with the reduced input dimensions.
While the aforementioned beliefs and concerns may appear intuitive and reasonable, there lacks strong empirical evidence confirming that standard BO is indeed incapable of high-dimensional optimization. To bridge this gap, we systematically investigated BO with standard GP regression across a variety of synthetic and real-world benchmark problems. Surprisingly, we found that, with ARD Matérn kernels and UCB as the acquisition function, the performance of standard BO consistently ranks among the best and often surpasses existing methods by a substantial margin. Those methods are specifically designed for high-dimensional optimization. This leads us to call for a re-evaluation and in-depth study of the potential of standard BO for high-dimensional optimization. The major discoveries are summarized as follows.
  • Optimization performance. We investigated BO with standard GP regression across eleven benchmarks widely used in literature, including six synthetic and five real-world high-dimensional optimization tasks. The number of optimization variables varies from 30 to 388. In comparison with six state-of-the-art high-dimensional BO methods, standard BO with ARD Matérn kernel and UCB consistently achieves the best (or close to the best) optimization performance across all the tasks. In contrast, the competing methods demonstrate strong performance when the target function aligns with their structural assumption. However, in cases of mismatch or when the structure of the target function is agnostic, such as those in real-world problems, the performance of the competing methods often experiences a significant drop. This underscores that BO with standard GP not only can excel in high-dimensional optimization but also is robust in accommodating various structures within the target functions.
  • Surrogate Learning. We investigated the task of learning high-dimensional functions with standard GP regression. Across ten benchmark datasets, our findings reveal that the prediction of standard GP equipped with Matérn kernels aligns well with the ground-truth, mostly performing better than more complex GP models that induce low-rank structures. This observation is in line with the recent results in kernel operator learning [3], where inputs are discretized functions — typically high-dimensional (hundreds or thousands of dimensions). It confirms that standard GP can serve as a capable surrogate for high-dimensional functions.
  • Computational Efficiency. We discovered that, with standard GP, achieving promising optimization performance is possible by using MAP estimation with diffuse priors or just maximum likelihood estimation. The marginalization of the parameters (kernel parameters and noise variance) via Markov-Chain Monte Carlo (MCMC) provides minor benefits while significantly increasing the training cost. In contrast, due to the usage of more complex surrogate models, training for alternative high-dimensional BO methods is often more expensive. Some of these methods heavily rely on extensive MCMC sampling to prevent overfitting, making the training of the surrogate model a computational bottleneck.
  • Failure Modes with Other Settings. We found failure modes of standard BO with other popular settings, such as using vanilla Square-Exponential (SE) kernels, Expected-Improvement (EI) and Thompson sampling (TS) as acquisition functions, and MAP estimation with concentrated Gamma priors over length-scale parameters. We conducted a careful ablation study to reveal the failure of these choices in both optimization and function learning, providing analysis and discussions. In Table 1, we list the discovered favorable and unfavorable settings for high-dimensional optimization.
@@ 2 Standard Bayesian Optimization
Consider maximizing a $d$-dimensional black-box objective function $f:\mathcal{X}\subset\mathbb{R}^{d}\rightarrow\mathbb{R}$. The form of $f(\cdot)$ is unknown. We can only query the function value at any input ${\bf x}\in\mathcal{X}$, possibly with noises, and no gradient information is accessible. We aim to find ${\bf x}^{\dagger}=\operatornamewithlimits{argmax}_{{\bf x}\in\mathcal{X}}f({\bf x})$. To achieve this, Bayesian Optimization (BO) employs a probabilistic surrogate model to learn and predict $f(\cdot)$ across the domain $\mathcal{X}$, while also quantifying the uncertainty of the prediction. This information is then integrated to compute an acquisition function, which measures the utility of querying at any new input location given the current function estimate. The utility typically encapsulates an exploration-exploitation trade-off. By maximizing the acquisition function, BO identifies a new input location at which to query, ideally closer to the optimum. Concurrently, the acquired new example is incorporated into the training dataset, and the surrogate model is retrained to improve the accuracy. BO begins by querying at a few (typically randomly selected) input locations, and trains an initial surrogate model. Then the iterative procedure repeats until BO converges to the optimum or a predefined stopping criterion is met.
The standard BO adopts GP regression [33] for surrogate modeling. Specifically, a GP prior is placed over the target function, $f\sim\mathcal{GP}\left(m({\bf x}),\kappa({\bf x},{\bf x}^{\prime})\right)$, where $m({\bf x})$ is the mean function, which is often set as a constant, and $\kappa(\cdot,\cdot)$ is the covariance function, which is often chosen as a Mercer kernel function. For example, a popular kernel used in BO is the Matérn-5/2 kernel with Automatic Relevance Determination (ARD),
@@
where $\tau=\sqrt{({\bf x}-{\bf x}^{\prime})^{\top}{\rm diag}(\frac{1}{\bm{\eta}})({\bf x}-{\bf x}^{\prime})}$, $a>0$ is the amplitude, and $\bm{\eta}>{\bf 0}$ are the length-scale parameters. We refer to this kernel as ARD because each input dimension has a distinct length-scale parameter. Alternatively, one can use a single length-scale parameter shared by all the input dimensions.
Suppose we have collected a set of training inputs ${\bf X}=[{\bf x}_{1},\ldots,{\bf x}_{N}]^{\top}$ and (noisy) outputs ${\bf y}=[y_{1},\ldots,y_{N}]^{\top}$. Due to the GP prior, the function values at ${\bf X}$, denoted as ${\bf f}=[f({\bf x}_{1}),\ldots,f({\bf x}_{N})]^{\top}$, follow a multi-variate Gaussian distribution, $p({\bf f})=\mathcal{N}({\bf f}|{\bf m},{\bf K})$, where ${\bf m}=[m({\bf x}_{1}),\ldots,m({\bf x}_{N})]^{\top}$, ${\bf K}$ is the kernel matrix on ${\bf X}$, and each $[{\bf K}]_{ij}=\kappa({\bf x}_{i},{\bf x}_{j})$. Next, a Gauss noise model is used to model the likelihood of ${\bf y}$, namely, $p({\bf y}|{\bf f})=\mathcal{N}({\bf y}|{\bf f},\sigma^{2}{\bf I})$ where $\sigma^{2}$ is the noise variance. One can marginalize ${\bf f}$ out to obtain
$p({\bf y}|{\bf X})=\mathcal{N}({\bf y}|{\bf m},{\bf K}+\sigma^{2}{\bf I}).$ (2)
To estimate the GP parameters (e.g., $a$, $\bm{\eta}$, and $\sigma^{2}$), one can place a prior over these parameters and use MCMC to compute posterior samples or perform MAP estimation. One can also simply maximize (2) to perform MLE. The predictive distribution (given the GP parameters) has a closed form: at any new input ${\bf x}^{*}$, the predictive distribution is conditional Gaussian,
$p\left(f({\bf x}^{*})|{\bf y}\right)=\mathcal{N}\left(f({\bf x}^{*})|\mu({\bf x}^{*}),v({\bf x}^{*})\right),$ (3)
where $\mu({\bf x}^{*})=m({\bf x}^{*})+\kappa({\bf x}^{*},{\bf X})({\bf K}+\sigma^{2}{\bf I})^{-1}({\bf y}-{\bf m})$, and $v({\bf x}^{*})=\kappa({\bf x}^{*},{\bf x}^{*})-\kappa({\bf x}^{*},{\bf X})({\bf K}+\sigma^{2}{\bf I})^{-1}\kappa({\bf X},{\bf x}^{*})$.
In each iteration, given the current GP surrogate, we maximize an acquisition function to identify the next input location at which to query. One commonly used acquisition function is the upper confidence bound (UCB), defined as
$\text{UCB}({\bf x})=\mu({\bf x})+\lambda\sqrt{v({\bf x})},\;\;{\bf x}\in\mathcal{X},$ (4)
where $\lambda$ represents the exploration level. There have been other popular acquisition functions, such as Expected Improvement (EI) [14], Thompson sampling (TS) [27], among others.
@@ 3 High Dimensional Bayesian Optimization
An enduring and widespread belief is that when the input dimension $d$ is high (e.g., a few hundreds), the standard BO is prone to failure. This belief might partly arise from an intuition that commonly used kernels, such as in (1), could encounter challenges in dealing with high-dimensional inputs, making GP struggle in capturing the target function. As a result, there is a dedicated line of research aimed at developing novel high-dimensional BO methods. The key idea of these methods is to introduce strong structural assumptions in surrogate modeling to sidestep directly handling the high-dimensional inputs in kernels.
3.1 Structural Assumption in Functional Space
The first class of methods assumes a decomposition structure within the functional space. Specifically, the target function $f$ is modeled as
@@ 3.1 Structural Assumption in Functional Space
where each ${\bf x}^{j}\subset{\bf x}$ is a small group of input variables, and ${\bf x}={\bf x}^{1}\cup\ldots\cup{\bf x}^{M}$. There can be a variety of choices for the group number $M$ and each variable group ${\bf x}^{j}$. In [16], the input ${\bf x}$ is partitioned into non-overlapped groups. After every a few BO iterations, a new partition is selected from a set of random partitions. The selection is based on the model evidence. In [26], the variable groups can overlap, and are represented as the maximum cliques on a dependency graph. A Gibbs sampling method was developed to learn the structure of the dependency graph. Han et al., [10] proposed Tree-UCB, which restricted the dependency graph to be tree-structured so as to boost the efficiency of structure learning. However, the most recent work [34] points out that, learning the group structure through relatively small data can be misleading. A wrongly learned structure can cause queries to be stuck at local optima. The work proposes RDUCB that randomly decomposed ${\bf x}$ into a tree structure, and surprisingly works better than those learned structures through various methods.
3.2 Structural Assumption in Input Space
Another class of methods assumes a low-rank structure in the input space.

### v4

@@ Abstract
A long-standing belief holds that Bayesian Optimization (BO) with standard Gaussian processes (GP) — referred to as standard BO — underperforms in high-dimensional optimization problems. While this belief seems plausible, it lacks both robust empirical evidence and theoretical justification. To address this gap, we present a systematic investigation. First, through a comprehensive evaluation across eleven widely used benchmarks, we found that while the popular Square Exponential (SE) kernel often leads to poor performance, using Matérn kernels enables standard BO to consistently achieve top-tier results, frequently surpassing methods specifically designed for high-dimensional optimization. Second, our theoretical analysis reveals that the SE kernel’s failure primarily stems from improper initialization of the length-scale parameters, which are commonly used in practice but can cause gradient vanishing in training. We provide a probabilistic bound to characterize this issue, showing that Matérn kernels are less susceptible and can robustly handle much higher dimensions. Third, we propose a simple robust initialization strategy that dramatically improves the performance of the SE kernel, bringing it close to state-of-the-art methods, without requiring any additional priors or regularization. We prove another probabilistic bound that demonstrates how the gradient vanishing issue can be effectively mitigated with our method. Our findings advocate for a re-evaluation of standard BO’s potential in high-dimensional settings.
@@ 1 Introduction
Despite numerous successes, there has been a widespread belief that BO with standard GP regression, referred to as standard BO, is limited to low-dimensional optimization problems [8, 23, 6, 22, 18, 31, 19]. It is commonly thought that the number of optimization variables should not exceed 15 or 20, as beyond this threshold, BO is prone to failure [8, 23]. A continuous line of research is dedicated to developing novel high-dimensional BO methods. The fundamental strategy is to impose strong structural assumptions into surrogate modeling so as to avoid directly dealing with high-dimensional inputs. One class of methods assumes a decomposition structure within the functional space [16, 24, 10, 33], where the target function is expressed as the summation of a group of low-dimensional functions, each operating over a small number of variables. Another family of methods assumes that the inputs are intrinsically low-rank. These methods either project the original inputs into a low-dimensional space [31, 23, 18] or use sparse-inducing priors to trim down massive input variables [6]. Subsequently, GP surrogates are built with the reduced input dimensions.
While the aforementioned concerns may seem plausible, there is a lack of both strong empirical evidence and theoretical justification to confirm and explain why standard BO would be ineffective in high-dimensional optimization. To bridge this gap, we systematically investigated standard BO in this paper. The major contributions of our work are summarized as follows:
  • Empirical Results. We investigated BO with standard GP across eleven widely-used benchmarks, including six synthetic and five real-world high-dimensional optimization tasks. The number of variables ranged from 30 to 388. We compared standard BO with nine state-of-the-art high-dimensional BO methods, and performed extensive evaluations and ablation studies. Surprisingly, while the popular SE kernel often led to poor performance, switching to the ARD Matérn kernel enabled standard BO to consistently achieve the best (or near-best) optimization performance across all the tasks. When successful, standard BO seems more flexible in accommodating various structures within the target functions.
  • Theory. Through analyzing the gradient structure of the GP likelihood, we identified the primary failure mode of standard BO, in particular when using the SE kernel. Specifically, the commonly used initializations for the length-scale parameters, such as setting them to one, is improper in high-dimensional settings and can easily cause gradient vanishing, preventing effective training. We proved a probabilistic tail bound to characterize this issue under mild conditions. Applying the same analytical framework to the Matérn kernel, we show through a comparison of the tail bounds that the Matérn kernel is less prone to gradient vanishing with the same length-scale initialization, making it more effective for handling high-dimensional problems.
  • Simple Robust Initialization. Based on our theoretical results, we proposed a simple yet robust length-scale initialization method, without requiring any additional priors or regularization. We proved a probabilistic bound showing that the probability of gradient vanishing decreases exponentially with the increase in dimensionality, implying that the gradient vanishing issue can be effectively mitigated in high-dimensional settings. Empirical evaluations demonstrate that our initialization method dramatically improves the performance of standard BO with SE kernels, enabling it to achieve, or come close to, the state-of-the-art performance in high-dimensional optimization.
@@ 2 Standard Bayesian Optimization
Consider maximizing a $d$-dimensional objective function $f:\mathcal{X}\subset\mathbb{R}^{d}\rightarrow\mathbb{R}$. The form of $f(\cdot)$ is unknown. We can query the function value at any input ${\bf x}\in\mathcal{X}$, possibly with noises, but no gradient information is accessible. We aim to find ${\bf x}^{\dagger}=\mathrm{argmax}_{{\bf x}\in\mathcal{X}}f({\bf x})$. To achieve this, Bayesian Optimization (BO) employs a probabilistic surrogate model to predict $f$ across the domain $\mathcal{X}$, while also quantifying the uncertainty of the prediction. This information is then integrated to compute an acquisition function, which measures the utility of querying at any new input location given the current function estimate. The utility typically encapsulates an exploration-exploitation trade-off. By maximizing the acquisition function, BO identifies a new input location at which to query, ideally closer to the optimum. Concurrently, the acquired new example is incorporated into the training dataset, and the surrogate model is retrained to improve the accuracy. BO begins by querying at a few (typically randomly selected) input locations, and trains an initial surrogate model. Then the iterative procedure repeats until convergence or a predefined stopping criterion is met.
The standard BO adopts GP regression [32] for surrogate modeling. A GP prior is placed over the target function, $f\sim\mathcal{GP}\left(m({\bf x}),\kappa({\bf x},{\bf x}^{\prime})\right)$, where $m({\bf x})$ is the mean function, which is often set as a constant, and $\kappa(\cdot,\cdot)$ is the covariance function, which is usually chosen as a Mercer kernel function. One most popular kernel used in BO is the Square Exponential (SE) kernel with Automatic Relevance Determination (ARD),
$\kappa_{\text{SE}}({\bf x},{\bf x}^{\prime})=a\exp(-\rho^{2}),$ (1)
where $a>0$ is the amplitude, $\rho=\sqrt{({\bf x}-{\bf x}^{\prime})^{\top}{\rm diag}(\frac{1}{\bm{\ell}^{2}})({\bf x}-{\bf x}^{\prime})}$, and $\bm{\ell}=(\ell_{1},\ldots,\ell_{d})^{\top}>{\bf 0}$ are the length-scale parameters. We refer to this kernel as ARD because each input dimension has a distinct length-scale parameter. An alternative choice is the ARD Matérn-5/2 kernel,
@@
Given training inputs ${\bf X}=[{\bf x}_{1},\ldots,{\bf x}_{N}]^{\top}$ and (noisy) outputs ${\bf y}=[y_{1},\ldots,y_{N}]^{\top}$, let us denote the function values at ${\bf X}$ as ${\bf f}=[f({\bf x}_{1}),\ldots,f({\bf x}_{N})]^{\top}$, which according to the GP prior, follow a multi-variate Gaussian distribution, $p({\bf f})=\mathcal{N}({\bf f}|{\bf m},{\bf K})$, where ${\bf m}=[m({\bf x}_{1}),\ldots,m({\bf x}_{N})]^{\top}$, ${\bf K}$ is the kernel matrix on ${\bf X}$, and each $[{\bf K}]_{ij}=\kappa({\bf x}_{i},{\bf x}_{j})$. Then one can employ a Gaussian likelihood for the observations ${\bf y}$, and obtain
$p({\bf y}|{\bf X})=\mathcal{N}({\bf y}|{\bf m},{\bf K}+\sigma^{2}{\bf I}),$ (3)
where $\sigma^{2}$ is the noise variance. To estimate the GP parameters, e.g., the length-scales $\bm{\ell}$, one can maximize the marginal likelihood or use MCMC sampling. The predictive distribution (given the GP parameters) at any new input ${\bf x}^{*}$, is a closed-form, conditional Gaussian distribution, $p\left(f({\bf x}^{*})|{\bf y}\right)=\mathcal{N}\left(f({\bf x}^{*})|\mu({\bf x}^{*}),v({\bf x}^{*})\right)$, where $\mu({\bf x}^{*})=m({\bf x}^{*})+\kappa({\bf x}^{*},{\bf X})({\bf K}+\sigma^{2}{\bf I})^{-1}({\bf y}-{\bf m})$, and $v({\bf x}^{*})=\kappa({\bf x}^{*},{\bf x}^{*})-\kappa({\bf x}^{*},{\bf X})({\bf K}+\sigma^{2}{\bf I})^{-1}\kappa({\bf X},{\bf x}^{*})$.
In each iteration, given the current GP surrogate model, an acquisition function is maximized to identify the next input location at which to query. One commonly used acquisition function is the upper confidence bound (UCB), defined as $\text{UCB}({\bf x})=\mu({\bf x})+\lambda\sqrt{v({\bf x})},\;\;{\bf x}\in\mathcal{X}$, where $\lambda$ represents the exploration level. There have been other popular acquisition functions, such as Expected Improvement (EI) [15], Thompson sampling (TS) [25], the recently proposed log-EI [1], among others.
@@ 3 High Dimensional Bayesian Optimization
An enduring and widespread belief is that when the input dimension $d$ is high, e.g., a few hundreds, the standard BO is prone to failure. This belief might partly arise from an intuition that commonly used kernels, such as (1), could encounter challenges in dealing with high-dimensional inputs, making GP struggle in capturing the target function. A dedicated line of research is developing novel high-dimensional BO methods. The key idea of these methods is to impose some structural assumptions in surrogate modeling to sidestep directly handling the high-dimensional inputs in kernels.
Structural Assumption in Functional Space. The first class of methods assumes a decomposition structure within the functional space. Specifically, the target function $f$ is modeled as
@@
where each ${\bf x}^{j}\subset{\bf x}$ is a small group of input variables, and ${\bf x}={\bf x}^{1}\cup\ldots\cup{\bf x}^{M}$. There can be a variety of choices for the group number $M$ and each variable group ${\bf x}^{j}$. In [16], the input ${\bf x}$ is partitioned into non-overlapped groups. After every a few BO iterations, a new partition is selected from a set of random partitions. The selection is based on the model evidence. In [24], the variable groups can overlap, and are represented as the maximum cliques on a dependency graph. A Gibbs sampling method was developed to learn the structure of the dependency graph. Han et al., [10] proposed Tree-UCB, which restricted the dependency graph to be tree-structured so as to boost the efficiency of structure learning. However, the more recent work [33] points out that, learning the group structure through relatively small data can be misleading. A wrongly learned structure can cause queries to be stuck at local optima. The work proposes RDUCB that randomly decomposes ${\bf x}$ into a tree structure, and surprisingly works better than those learned structures through various methods.